OpenStudy (anonymous):

find critical numbers of f(x) = x cubed + x squared - x... can someone help me set the derivative to zero i know f'(x) = 3x^2 +2x - 1 i do not know what to do with the "-1" in factoring and cant find an example. find critical numbers of f(x) = x cubed + x squared - x... can someone help me set the derivative to zero i know f'(x) = 3x^2 +2x - 1 i do not know what to do with the "-1" in factoring and cant find an example. @Mathematics

OpenStudy (anonymous):

is the correct answer -1 for smaller value and 1/3 for maximum?

OpenStudy (amistre64):

working with calculus can make you forget your algebra

OpenStudy (hoblos):

3x^2 +2x-1 is a quadratic of the form ax^2 +bx+c but if a+c=b=> the solution are x=-1 and x=-c/a so the solutions for your quadratic are x=-1 and x=1/3

OpenStudy (amistre64):

(3x-1)(x+1) is the factoring if i did it in me head right